OpenStudy (anonymous):

Find the exact distance between the points (7,-22)(-8,17)

OpenStudy (shane_b):

Use the distance formula:\[d=\sqrt{(x_2-x_1)^2)+(y_2-y_1)^2)}\]

OpenStudy (anonymous):

would d=1,551?

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

did you use the formula @Shane_B gave you?

OpenStudy (anonymous):

Yes. d=1,746?

OpenStudy (helder_edwin):

if P(7,-22) and Q(-8,17) then \[ \LARGE{ d(P,Q)=\sqrt{(7-(-8))^2+(-22-17)^2}=\\ \sqrt{15^2+39^2}=\sqrt{1746}=... }\]

OpenStudy (anonymous):



OpenStudy (helder_edwin):

yes but 194 is not prime so you can find more of its prime factors

OpenStudy (helder_edwin):

sorry you did it great

OpenStudy (anonymous):

thanks so much for your help....so leave the 194 as is?